OpenStudy (anonymous):

What represents 4x^5/6 in radical form?

ganeshie8 (ganeshie8):

hint : \[\huge a^{\frac{m}{n}} = \sqrt[n]{m}\]

ganeshie8 (ganeshie8):

\[\huge x^{\frac{5}{6}} = ?\]

OpenStudy (anonymous):

Sorry my computer crashed its 6sqrx^5

ganeshie8 (ganeshie8):

yes : \[\huge 4x^{\frac{5}{6}} =4\sqrt[6]{x^5} \]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I understand now

ganeshie8 (ganeshie8):

good to hear, yw!